Suppose ABC manufacturer has the following costs and sales expectations



 Variable cost per units(unit variable cost)……….10br



 Total Fixed cost……………...300,000 br



 Expected unit sales…………..50,000units



 Total invested capital ………..1milion br



And assume the ABC manufacturer wants to earn a 20 percent markup on sales and 20 percent



return on investment.



a. What should be the price of ABC manufacturer product and the profit per unit



when it uses markup pricing method?



b. What should be the price of ABC manufacturer product and the profit per unit



when it uses investment return pricing method?



c. Compute the break-even volume�

1
Expert's answer
2022-02-25T08:39:01-0500

A

Variable cost = 10*50,000 = 500,000

Fixed cost                             =300,000

Total cost                              =800,000

Mark up (20*800,000)        =160,000

                                                =960,000

Price = 960,000 / 50,000

         =19.2 

 

 

 

B

Variable cost =                        500,000

Fixed cost       =                       300,000

Total cost        =                        800,000

ROI          =20*1000, 000=         200,000

 

Price = 1000,000 / 50,000

                =20

 

 

C

 

BEP= fixed cost /contribution unit

Contribution analysis = selling price -    variable cost

Mark up = 19.2 – 10 = 9.2

ROI  = 20 - 10 = 10

BEP-mark up=300,000/9.2 =  326087 units

ROI = 300,000/ 10  =30,000 units